The Prince-Robbins partnership has the following capital account balances on January 1, 2015:

  

  Prince, Capital $ 145,000
  Robbins, Capital 135,000

  

      Prince is allocated 80 percent of all profits and losses with the remaining 20 percent assigned to Robbins after interest of 10 percent is given to each partner based on beginning capital balances.

      On January 2, 2015, Jeffrey invests $82,000 cash for a 20 percent interest in the partnership. This transaction is recorded by the goodwill method. After this transaction, 10 percent interest is still to go to each partner. Profits and losses will then be split as follows: Prince (50%), Robbins (30%), and Jeffrey (20%). In 2015, the partnership reports a net income of $35,000.

  

a.

Prepare the journal entry to record Jeffrey entrance into the partnership on January 2, 2015. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

Record the entry for goodwill allocation, during the admission of a new partner.

Record the cash received from new partner.

b.

Determine the allocation of income at the end of 2015.

Income Allocation
Prince
Robbins
Jeffery
Answer #1

Requirement: Calculation of Goodwill: Total Capital Actual Capital Goodwill =82000/20% =145000+135000+82000 $ $ $ 4,10,000 3,